OpenStudy (lacris):

Please help me, I will give you a shiny medal :) 1. Suppose that a single card is selected from a standard​ 52-card deck. What is the probability that the card drawn is a club? 2. Now suppose that a single card is drawn from a standard​ 52-card deck, but it is told that the card is black. What is the probability that the card drawn is a club​? The probability that the card drawn from a standard​ 52-card deck is a club is ??

OpenStudy (lacris):

For the first question, I did 13/52 = 0.25

OpenStudy (lacris):

Because there can be 13 clubs, spades, hearts, and diamonds. However with the second question, I'm not sure, because of the conditional probability rule..

OpenStudy (lacris):

oh, never mind I actually got it correct, because the next question is asking for two of the black section of cards so: 13+13 = 26/52 = 0.5 which the probability.